OpenStudy (anonymous):

Utako sells environmentally friendly cleaning products. Her income consists of $200 per week plus a commission of 12% of her sales. Write an algebraic expression for Utako's weekly income, I, in terms of her sales, S. [Note: Variables are case sensitive:"I" and "i" are not the same!]

OpenStudy (stamp):

weekly income I = 200 $ + commission commission = 12% sales = 12% S \[I = 200 + 12\% S\]\[I = 200 + \frac{ 12 }{ 100 }S\]